What information does the payback period provide? Suppose Extensive Enterprisess CFO is evaluating a project with the followPlease help!Thank you!
Answer #1

1:453531

2: Option 1

Since payback is 2.5 years

Initial cost = CF1+CF2+0.5*CF3 = 300000+400000+0.5*450000

=925000

NPV= C0+ CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

=-925000+300000/1.07^1+400000/1.07^2+450000/1.07^3+500000/1.07^4

=453531

Option 1: The discounted payback will only consider the cash flows till the time the initial outlay is recovered.

Option 2 and 3 are not true. The discounted payback considers cash flows and takes into account the discounted cash flows.